how do you solve : 2+7n=8+n?

Found 2 solutions by solver91311, annika21:
Answer by solver91311(24713)



Add to both sides.

Add to both sides.

Multiply both sides by


John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by annika21(5)

how do you solve : 2+7n=8+n?

7n-1n=8-2 (brought everything to each side)
6n=6
n=1 (Divide